maria purchased 12 pens and 8 notebooks for $15.00. bob bought 4 pens and 3 notebooks for $5.50. find the price of one pen and the price of one notebook ,agebraically.
Answer by checkley75(3666)

12P+8N=15.00
4P+3N=5.50 NOW MULTIPLY THE SECOND EQUATION BY -3 & ADD:
-12P-9N=-16.50
12P+8N=15.00
-----------------------------
-N=-1.50
N=1.50 PRICE OF THE NOTEBOOK.
12P+8*1.50=15.00
12P+12.00=1500
12P=15.00-12.00
12P=3.00
P=3.00/12
P=.25 COST OF THE PEN.
PROOF
4*.25+3*1.50=5.50
1.00+4.50=5.50
5.50=5.50